A motorcyclist is making an electric vest that, when connected to the motorcycle's 12 V battery, will warm her on cold rides. She is using 0.23-mm-diameter copper wire, and she wants a current of 5.1 A in the wire.What length wire must she use?
Answer #1

V = IR

12 = 5.1*R

=> R = 12/5.1 = 2.352 ohm

R = resistivity*L/A


=> 2.352 = 1.68*10^-8*L/(3.14*(0.23/2*10^-3)^2)

=> L =   (2.352*(3.14*((0.23/2)*10^-3)^2))/(1.68*10^-8) = 5.81 m